# Approximating the coefficients of the Bessel functions

### v3

@@ 10.5. Bessel generating functions for exponentially decaying probability measures
(A): We can follow the proof of Corollary˜1.7 that involves first proving convergence in terms of moments and using this fact to justify convergence in distribution. Instead of using Theorem˜1.1, we use Corollary˜10.18.
(B): We use Lemma˜10.19 to deduce that for $N\geq 1$, $\mu^{N}$ is compactly supported and therefore exponentially decaying. Afterwards, we follow the proof of (A) for the remaining steps of the proof. ∎
We can also state the analogues of Corollary˜10.20 for the type BC and D root systems, which are similar to Corollary˜1.8 and part (B) of Corollary˜1.7, respectively. To determine what the differences are, it suffices to compare Corollary˜10.20 with part (A) of Corollary˜1.7.
@@
(36) $G_{\mu_{t}(a)}^{A^{N-1}(\theta)}(x)=B_{a}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
@@
Assume that $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$. Then,
$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=G_{\eta}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
over $\overline{B}_{R}(0)$ for all $t\geq 0$.
@@
See Lemma˜10.16 and (36). ∎
@@
Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.22 implies that
@@
Assume that $\alpha>0$. We can follow the method of the proof of Corollary˜10.20. First, since condition (b) of Corollary˜10.18 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well. The next step is to deduce that $\mu_{\alpha N}(\eta_{N})$ is exponentially decaying for all $N\geq 1$. This does not follow from (37) and Lemma˜10.19 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
Suppose $y\in\mathbb{R}^{N}$. From [airy_beta]*(23), we have that the density of $y$ for $DBM(\theta)$ at time $\alpha N$ conditioned on the initial value $a$ is at most
$\exp\left(-\frac{(\left\lVert a\right\rVert_{2}-\left\lVert y\right\rVert_{2})^{2}}{2\alpha N}\right)y^{O(1)},$
where we have applied the upper bound $B_{a}\left(\frac{y}{\alpha N}\right)\leq\sqrt{N!}\exp\left(\frac{\left\lVert a\right\rVert_{2}\left\lVert y\right\rVert_{2}}{\alpha N}\right)$ that we deduce from part (e) of Lemma˜2.6.
Assume that $\eta_{N}$ exponentially decays at rate $R$. Then, we have that for all $\epsilon>0$,
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2\alpha N}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
and
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta_{N},\\ y\sim\mu_{\alpha N}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
This implies that $\mu_{\alpha N}(\eta_{N})$ exponentially decays at any rate less than $R$.
Then, because
$G_{\mu_{\alpha N}(\eta_{N})}^{A^{N-1}(\theta)}(x)=G_{\eta_{N}}^{A^{N-1}(\theta)}(x)\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
over $\overline{B}_{R}(0)$ by Lemma˜10.22, applying Corollary˜10.18 again gives that the $k$th free cumulant of $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$. Hence, the probability measures $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to the free convolution of $\eta$ and the semicircular law multiplied by $\frac{\alpha}{\theta}$ in terms of moments as $N\rightarrow\infty$. If $\eta$ is compactly supported, then the free convolution is compactly supported and therefore uniquely determined by its moments by Carleman’s condition, which implies that the sequence of probability measures also converges weakly to the free convolution after applying [billingsley]*Theorem 30.2.

### v4

@@ 10.5. Bessel generating functions for exponentially decaying probability measures
(A): We can follow the proof of Corollary˜1.7 that involves first proving convergence in terms of moments and using this fact to justify convergence in distribution. Instead of using Theorem˜1.1, we use Corollary˜10.17.
(B): We use Lemma˜10.21 to deduce that for $N\geq 1$, $\mu^{N}$ is compactly supported and therefore exponentially decaying. Afterwards, we follow the proof of (A) for the remaining steps of the proof. ∎
In the following two corollaries, we use notation defined in Corollary˜1.8.
Corollary 10.23.
[…]
  1. (A)
    Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are exponentially decaying Borel probability measures over $\mathbb{C}^{N}$ for $N\geq 1$ such that:
    1. 1.
      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
    2. 2.
      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta_{0}N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
    For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
  2. (B)
    Suppose $\theta_{0}$ and $\theta_{1}$ are nonnegative. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true. For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure such that $G_{\mu^{N}}^{BC^{N}(\theta_{0},\theta_{1})}=G_{\mu_{a}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}G_{\mu_{b}^{N}}^{BC^{N}(\theta_{0},\theta_{1})}$ in a neighborhood of the origin.
    Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]$ converges to the rectangular free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ with $\lambda$ set as $\frac{1}{1+c}$ in distribution and in terms of moments.
Remark 10.24.
Note that in part (A) of the previous corollary, the rectangular free convolution is not defined in [rectangular_free_convolution] when $c$ is complex. In this case, we define it using part (B) of Corollary˜10.2.
Corollary 10.25.
  1. (A)
[…]
    1. 1.
      The measure $\operatorname{\mathbb{E}}_{a\sim\mu_{a}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{a_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{a}$ over $\mathbb{C}$ in terms of moments.
    2. 2.
      The measure $\operatorname{\mathbb{E}}_{b\sim\mu_{b}^{N}}\left[\sum_{i=1}^{N}\frac{1}{N}\delta\left(\frac{b_{i}}{\theta N}\right)\right]$ converges to the compactly supported Borel probability measure $\mu_{b}$ over $\mathbb{C}$ in terms of moments.
    For $N\geq 1$, suppose $\mu^{N}$ is an exponentially decaying Borel probability measure such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
  2. (B)
    Suppose $\theta\in\mathbb{C}$ is positive. Suppose $\mu_{a}^{N}$ and $\mu_{b}^{N}$ are compactly supported Borel probability measures over $\mathbb{R}^{N}$ for $N\geq 1$ and $\mu_{a}$ and $\mu_{b}$ are compactly supported Borel probability measures over $\mathbb{R}$ such that conditions 1 and 2 from (A) are true. For $N\geq 1$, suppose $\mu^{N}$ is a Borel probability measure over $\mathbb{R}^{N}$ such that $G_{\mu^{N}}^{D^{N}(\theta)}=G_{\mu_{a}^{N}}^{D^{N}(\theta)}G_{\mu_{b}^{N}}^{D^{N}(\theta)}$ in a neighborhood of the origin. Then, the measure $\mathbb{E}_{a\sim\mu^{N}}\left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta N}\right)\right]$ converges to the free convolution of $\frac{1}{2}(\mu_{a}+\mu_{a}^{-})$ and $\frac{1}{2}(\mu_{b}+\mu_{b}^{-})$ in distribution and in terms of moments.
@@
(36) $G_{\mu_{t}(a)}^{A^{N-1}(\theta)}(x)=J_{a}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)$
@@
Assume that $\eta$ is a Borel probability measure over $\mathbb{R}^{N}$ that exponentially decays at rate $R>0$. Then, for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$ and
$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x)$
over $B_{R}(0)$.
@@
First, we show that for all $t\geq 0$, $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$. This is obviously true if $t=0$, so assume that $t>0$.
Suppose $y\in\mathbb{R}^{N}$. From [airy_beta]*(23), we have that the density of $y$ for $DBM(\theta)$ at time $t$ conditioned on the initial value $a$ is at most
$\exp\left(-\frac{(\left\lVert a\right\rVert_{2}-\left\lVert y\right\rVert_{2})^{2}}{2t}\right)y^{O(1)},$
where we have applied the upper bound $B_{a}\left(\frac{y}{\alpha N}\right)\leq\sqrt{N!}\exp\left(\frac{\left\lVert a\right\rVert_{2}\left\lVert y\right\rVert_{2}}{\alpha N}\right)$ that we deduce from part (e) of Lemma˜2.6. Then, we have that for all $\epsilon>0$,
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}\geq(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$\leq\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(-\frac{((1-\frac{1}{1+\epsilon})\left\lVert y\right\rVert_{2})^{2}}{2t}+\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)y^{O(1)}\right]<\infty$
and
$\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp\left(\frac{R}{1+\epsilon}\left\lVert y\right\rVert_{2}\right)\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]$
$<\operatorname{\mathbb{E}}_{\begin{subarray}{c}a\sim\eta,\\ y\sim\mu_{t}(a)\end{subarray}}\left[\exp(R\left\lVert a\right\rVert_{2})\mathbf{1}\{\left\lVert y\right\rVert_{2}<(1+\epsilon)\left\lVert a\right\rVert_{2}\}\right]<\infty.$
This implies that $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$.
For the computation of the Bessel generating function $G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}$, we have that
$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{2N}}J_{a}^{A^{N-1}(\theta)}(x)d\mu_{t}(\eta)(a).$
for $x\in\mathbb{C}^{N}$. Because $\mu_{t}(\eta)$ exponentially decays at any rate less than $R$, we can use part (e) of Lemma˜2.6 and Fubini’s theorem to deduce that for $x\in B_{R}(0)$,
$G_{\mu_{t}(\eta)}^{A^{N-1}(\theta)}(x)=\int_{\mathbb{R}^{N}}\left(\int_{\mathbb{R}^{N}}J_{a(t)}^{A^{N-1}(\theta)}(x)d\mu_{t}(a)(a(t))\right)d\eta(a).$
Afterwards, applying (36) and part (A) of Lemma˜10.15 implies that this expression equals
$\int_{\mathbb{R}^{N}}J_{a}^{A^{N-1}(\theta)}(x)\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)d\eta(a)=\exp\left(\frac{t}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta}^{A^{N-1}(\theta)}(x).$
@@
Observe that $H_{\theta}^{t}$ clearly exponentially decays at any rate. Correspondingly, the measure’s Bessel generating function converges over $\mathbb{C}^{N}$. Furthermore, Lemma˜10.27 implies that
@@
Assume that $\alpha>0$. We can follow the method of the proof of Corollary˜10.22. First, since condition (b) of Corollary˜10.17 is satisfied when we set $\mu_{N}$ to be $\eta_{N}$ and $c_{(k)}$ to be the $k$th free cumulant of $\eta$ for all $k\geq 1$, we have that condition (a) is satisfied as well. The next step is to deduce that $\mu_{\alpha N}(\eta_{N})$ is exponentially decaying for all $N\geq 1$. This does not follow from (37) and Lemma˜10.21 because $H_{\theta}^{\sqrt{\alpha\theta N}}$ is not compactly supported.
Assume that $\eta_{N}$ exponentially decays at rate $R$. By Lemma˜10.27, $\mu_{\alpha N}(\eta_{N})$ exponentially decays at any rate less than $R$. Then, because
$G_{\mu_{\alpha N}(\eta_{N})}^{A^{N-1}(\theta)}(x)=\exp\left(\frac{\alpha N}{2}\sum_{i=1}^{N}x_{i}^{2}\right)G_{\eta_{N}}^{A^{N-1}(\theta)}(x)$
over $B_{R}(0)$ by Lemma˜10.27, applying Corollary˜10.17 again gives that the $k$th free cumulant of $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to $c_{(k)}+\mathbf{1}\{k=2\}\frac{\alpha}{\theta}$ as $N\rightarrow\infty$. Hence, the probability measures $\operatorname{\mathbb{E}}_{a\sim\mu_{\alpha N}(\eta_{N})}\left[\sum_{i=1}^{N}\frac{a_{i}}{\theta N}\right]$ converges to the free convolution of $\eta$ and the semicircular law multiplied by $\frac{\alpha}{\theta}$ in terms of moments as $N\rightarrow\infty$. If $\eta$ is compactly supported, then the free convolution is compactly supported and therefore uniquely determined by its moments by Carleman’s condition, which implies that the sequence of probability measures also converges weakly to the free convolution after applying [billingsley]*Theorem 30.2.